# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ Appendix D Rotating $\mathrm{SL}(2,\mathbb{R})$ drives
Our starting point is the Hamiltonian (5.3) expressed as
$H(t)=\frac{2\pi k}{L}\biggl[\cosh(2\lambda)K_{0}+\frac{i}{2}\sinh(2\lambda)\biggl(e^{ik\omega t}K_{-}-e^{-ik\omega t}K_{+}\biggr)\biggr]-\frac{\pi c}{12L}k^{2}$ (D.1)
@@
$K_{0}\equiv\frac{1}{k}\Bigl[L_{0}+\frac{c}{24}(k^{2}-1)\Bigr],\qquad K_{-}\equiv\frac{i}{k}L_{k},\qquad K_{+}\equiv-\frac{i}{k}L_{-k}.$ (D.2)
@@
$\left[K_{-},K_{+}\right]=2K_{0},\qquad\left[K_{0},K_{\pm}\right]=\pm K_{\pm},\qquad K_{0}^{\dagger}=K_{0},\qquad K_{-}^{\dagger}=K_{+}.$ (D.3)
By convention, the generators $K_{\pm}$ have a factor $i$, meaning that the corresponding coefficients $\beta_{\pm}=\mp\frac{i\pi k}{L}\sinh(2\lambda)e^{\mp ik\omega t}$ in Eq. (D.1) satisfy $\overline{\beta_{+}}=\beta_{-}$.
To make analytical progress, we use the fundamental, nonunitary representation of $\mathfrak{su}(1,1)$ in terms of $2\times 2$-matrices:
$K_{0}\big|_{2\times 2}=\begin{pmatrix}-1/2&0\\ 0&1/2\end{pmatrix},\qquad K_{-}\big|_{2\times 2}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\qquad K_{+}\big|_{2\times 2}=\begin{pmatrix}0&0\\ -1&0\end{pmatrix}$. (D.4)
The Hamiltonian (D.1) in this picture reads
$H(t)\big|_{2\times 2}=\frac{\pi k}{L}\begin{pmatrix}-\cosh(2\lambda)-ck/12&i\sinh(2\lambda)e^{ik\omega t}\\ i\sinh(2\lambda)e^{-ik\omega t}&\cosh(2\lambda)-ck/12\end{pmatrix}$. (D.5)
Naturally, this finite-dimensional representation is nonunitary since $\mathfrak{sl}(1,1)$ is noncompact—manifested already in Eq. (D.5) as the matrix is not Hermitian. However, as Eq. (D.4) is the fundamental representation of $\mathfrak{su}(1,1)$, it is faithful, so any lesson learned for this representation holds in general for $\mathfrak{su}(1,1)$.
@@
By definition of time-ordered exponentials as product integrals, the time-evolution operator (5.4) can be represented as in Eq. (5.5):
$U(t)\big|_{2\times 2}=e^{i\frac{\pi c}{12L}k^{2}t}\lim_{M\to\infty}\prod_{m=0,\ldots,M}^{\longleftarrow}\exp\biggl[\frac{\pi k}{L}\delta t\begin{pmatrix}i\cosh(2\lambda)&\sinh(2\lambda)e^{ik\omega m\delta t}\\ \sinh(2\lambda)e^{-ik\omega m\delta t}&-i\cosh(2\lambda)\end{pmatrix}\biggr$], (D.6)
@@
$\begin{pmatrix}0&e^{-i\omega s/2}\\ e^{i\omega s/2}&0\end{pmatrix}\begin{pmatrix}-i\bar{\alpha}&\bar{\beta}\\ \beta&i\alpha\end{pmatrix}\begin{pmatrix}0&e^{-i\omega s/2}\\ e^{i\omega s/2}&0\end{pmatrix}=\begin{pmatrix}i\alpha&\beta\,e^{-i\omega s}\\ \bar{\beta}\,e^{i\omega t}&-i\bar{\alpha}\end{pmatrix}$ (D.7)
for $s=-km\delta t$ and $\alpha=\cosh(2\lambda)$, $\beta=\sinh(2\lambda)$ to rewrite each factor in Eq. (D.6). As a result, each pair of consecutive factors gives rise to products of the form
$\begin{pmatrix}0&e^{ik\omega(m+1)\delta t/2}\\ e^{-ik\omega(m+1)\delta t/2}&0\end{pmatrix}\begin{pmatrix}0&e^{ik\omega m\delta t/2}\\ e^{-ik\omega m\delta t/2}&0\end{pmatrix}=\begin{pmatrix}e^{ik\omega\delta t/2}&0\\ 0&e^{ik\omega\delta t/2}\end{pmatrix}$, (D.8)
which thus are independent of $m$. The latter observation is crucial, since it implies that
$U(t)\big|_{2\times 2}=e^{i\frac{\pi c}{12L}k^{2}t}\lim_{M\to\infty}\begin{pmatrix}0&e^{ik\omega t/2}\\ e^{-ik\omega t/2}&0\end{pmatrix}\!\mathsf{A}^{M}\exp\biggl[\frac{\pi k}{L}\delta t\begin{pmatrix}-i\cosh(2\lambda)&\sinh(2\lambda)\\ \sinh(2\lambda)&i\cosh(2\lambda)\end{pmatrix}\biggr]\!\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, (D.9)
@@
$\mathsf{A}=\exp\biggl[\frac{\pi k}{L}\delta t\begin{pmatrix}-i\cosh(2\lambda)&\sinh(2\lambda)\\ \sinh(2\lambda)&i\cosh(2\lambda)\end{pmatrix}\biggr]\begin{pmatrix}e^{ik\omega\delta t/2}&0\\ 0&-e^{ik\omega\delta t/2}\end{pmatrix}=\begin{pmatrix}a&b\\ \bar{b}&\bar{a}\end{pmatrix}+O(\delta t^{2})$ (D.10)
with $a=\bigl[1-i(\pi k/L)\delta t\cosh(2\lambda)\bigr]e^{ik\omega\delta t/2}$ and $b=(\pi k/L)\delta t\sinh(2\lambda)e^{-ik\omega\delta t/2}$. Diagonalizing $\mathsf{A}$, noting that all $O(\delta t^{2})$ contributions give rise to subleading corrections in $1/M$, yields
$U(t)\big|_{2\times 2}$ $=e^{i\frac{\pi c}{12L}k^{2}t}\begin{pmatrix}e^{ik\omega t/2}&0\\ 0&e^{-ik\omega t/2}\end{pmatrix}$ (D.11)
@@
Now exploiting the fundamental representation (D.4) and the ansatz $e^{i[G+\bar{F}L_{-k}+EL_{0}+FL_{k}]}\big|_{2\times 2}$ ($G,E\in\mathbb{R}$, $F\in\mathbb{C}$) for the second matrix in Eq. (D.11), one can show that Eq. (5.6) holds for the evolution operator. This can also be used to find the diffeomorphism $g_{t}\in\operatorname{Diff}(S^{1})$ corresponding to the two last exponentials in Eq. (5.6). Owing to the properties of one-parameter flows, one can indeed show that $U(t)=e^{i\alpha_{t}}\,{\cal U}(g_{t})\,{\cal U}(g_{0})^{-1}$ with $g_{t}$ of the form (4.14), where $\zeta(\varphi)$ is a transformation (5.1) with $\alpha$, $\beta$ as in Eq. (5.7) and $\Omega$ is given by Eq. (5.8).
@@
When the initial condition is $g_{0}=f_{0}$, we also need to express ${\cal U}(f_{0})$ appearing in the initial state $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$ for $f_{0}$ of the $\mathrm{SL}(2,\mathbb{R})$ form (5.1) with $\alpha=\cosh(\lambda)$, $\beta=\sinh(\lambda)$ as a $2\times 2$ matrix. Since $f_{0}$ is the one-parameter flow generated by $(2/k)\sin(k\varphi)\partial_{\varphi}$ with respect to $\lambda$, one finds
${\cal U}(f_{0})\big|_{2\times 2}=\exp\biggl(\frac{\lambda}{k}[L_{-k}-L_{k}]\biggr)\big|_{2\times 2}=\begin{pmatrix}\cosh(\lambda)&i\sinh(\lambda)\\ -i\sinh(\lambda)&\cosh(\lambda)\\ \end{pmatrix}$. (D.12)
Replicating the steps that led to Eq. (5.6), the above can be used to factorize
$U(t)\,{\cal U}(f_{0})=e^{i\rho_{t}}e^{\chi_{t}L_{-k}-\bar{\chi}_{t}L_{k}}e^{i\sigma_{t}L_{0}},$ (D.13)
where
$\rho_{t}=\frac{\pi c}{12L}\Bigl[k^{2}t+\frac{L\sigma_{t}}{2\pi}\bigl(k^{2}-1\bigr)\Bigr],\qquad\sigma_{t}=-\omega t-\frac{2}{k}\arctan\Bigl(\Bigl[\frac{2\pi}{L\Omega}-\frac{\omega}{\Omega}\Bigr]\tan(k\Omega t/2)\Bigr),$ (D.14)
and $\eta_{t}$ in (5.10), with $\Omega$ given in Eq. (5.8). Evaluating $U(t)\,{\cal U}(f_{0})|h\rangle$ yields Eq. (5.9), after factoring out an overall phase that is equal to $\rho_{t}+h\sigma_{t}$ due to Eq. (2.16).
Lastly, combining Eqs. (D.11) and (D.12) allows us to represent ${\cal U}(f_{0})^{-1}U(t)\,{\cal U}(f_{0})$ as an $\mathrm{SU(1,1)}$ matrix:
${\cal U}(f_{0})^{-1}U(t)\,{\cal U}(f_{0})\big|_{2\times 2}=\begin{pmatrix}u_{t}&w_{t}\\ \bar{w_{t}}&\bar{u_{t}}\end{pmatrix}$ (D.15)
with $u_{t}$ as in Eq. (5.14) and $w_{t}=\sinh(2\lambda)\bigl[\frac{\omega}{\Omega}\sin(k\Omega t/2)\cos(k\omega t/2)-\cos(k\Omega t/2)\sin(k\omega t/2)+i\frac{2\pi}{L\Omega}\sin(k\Omega t/2)\sin(k\omega t/2)\bigr]$. On general grounds, this must be the nonunitary $2\times 2$-matrix representation of an element of the form $e^{D}e^{AL_{-k}}e^{BL_{0}}e^{CL_{k}}$, which implies
$A=\frac{i}{k}\frac{\bar{w}}{u},\qquad B=-\frac{2}{k}\log(u),\qquad C=\frac{i}{k}\frac{w}{u},\qquad e^{D}=e^{i\frac{\pi c}{12L}k^{2}t}e^{B\frac{c}{24}(k^{2}-1)}.$ (D.16)
It follows using Eq. (2.16) that
$\langle\psi(0)|\psi(t)\rangle=\langle h|\,{\cal U}(f_{0})^{-1}U(t)\,{\cal U}(f_{0})|h\rangle=e^{D}\langle h|e^{BL_{0}}|h\rangle=e^{D+Bh},$ (D.17)
which yields Eq. (5.13).
@@ Appendix E Lattice calculations
Lattice Hamiltonian.
We consider a driven inhomogeneous free-fermion chain at half-filling. Given fermionic creation (annihilation) operators $c_{j}^{\dagger}$ ($c_{j}$) acting on sites $j=1,\ldots,N$ and satisfying the usual anticommutation relations $\{c_{j^{\vphantom{\prime}}}^{\vphantom{\dagger}},c_{j^{\prime}}^{\dagger}\}=\delta_{j,j^{\prime}}$ and $\{c_{j^{\vphantom{\prime}}}^{\vphantom{\dagger}},c_{j^{\prime}}^{\vphantom{\dagger}}\}=\{c_{j^{\vphantom{\prime}}}^{\dagger},c_{j^{\prime}}^{\dagger}\}=0$, the Hamiltonian is
$H_{N}(t)=-\frac{1}{2}\sum_{j=1}^{N}J_{j}^{\vphantom{\dagger}}(t)\Bigl(c_{j}^{\dagger}c_{j+1}^{\vphantom{\dagger}}+\textrm{h.c.}\Bigr)$ (E.1)
with inhomogeneous nearest-neighbor couplings $J_{j}(t)=J_{j}(t+T)>0$ and periodic boundary conditions, $c_{N+1}^{(\dagger)}=c_{1}^{(\dagger)}$ and $J_{N+1}^{\vphantom{\dagger}}(t)=J_{1}^{\vphantom{\dagger}}(t)$. This model can be mapped (up to boundary terms) to the inhomogeneous spin chain in Eq. (1.2) with $\Delta=0$ via a Jordan-Wigner transformation. Assuming that the system size $L$ is much larger than the lattice spacing $L/N$, i.e. $N\gg 1$, and that $J_{j}(t)$ vary on mesoscopic length scales, the model (E.1) can be effectively described in the low-energy regime by an inhomogeneous CFT (1.1) with $c=1$ but involving both the right- and left-moving (chiral and antichiral) components of the CFT [26].
@@
In our lattice calculations, we only consider the fermions at zero temperature, meaning that our CFT results should be a good approximation of the lattice dynamics. Concretely, we let $|\psi_{N}(0)\rangle$ be the ground state of $H_{N}(0)$, which in the CFT language corresponds to a coherent state in the vacuum superselection sector ($h=\bar{h}=0$), and calculate the return probability

### v2

@@ Appendix D Rotating $\mathrm{\bf SL}\bm{(}\bm{2}\bm{,}\mathbb{R}\bm{)}$ drives
Our starting point is the Hamiltonian (73) expressed as
$H(t)=\frac{2\pi k}{L}\biggl[\cosh(2\lambda)K_{0}+\frac{i}{2}\sinh(2\lambda)\biggl(e^{ik\omega t}K_{-}-e^{-ik\omega t}K_{+}\biggr)\biggr]-\frac{\pi c}{12L}k^{2}$ (104)
@@
$K_{0}\equiv\frac{1}{k}\Bigl[L_{0}+\frac{c}{24}(k^{2}-1)\Bigr],\qquad K_{\mp}\equiv\pm\frac{i}{k}L_{\pm k}.$ (105)
@@
$\left[K_{-},K_{+}\right]=2K_{0},\qquad\left[K_{0},K_{\pm}\right]=\pm K_{\pm},\qquad K_{0}^{\dagger}=K_{0},\qquad K_{-}^{\dagger}=K_{+}.$ (106)
By convention, the generators $K_{\pm}$ have a factor $i$, meaning that the corresponding coefficients $\beta_{\pm}=\mp\frac{i\pi k}{L}\sinh(2\lambda)e^{\mp ik\omega t}$ in Eq. (104) satisfy $\overline{\beta_{+}}=\beta_{-}$.
To make analytical progress, we use the fundamental representation of $\mathfrak{su}(1,1)$ in terms of $2\times 2$-matrices:
[…]
The Hamiltonian (104) in this picture reads
$H(t)\big|_{2\times 2}=\frac{\pi k}{L}\begin{pmatrix}-\cosh(2\lambda)-ck/12&i\sinh(2\lambda)e^{ik\omega t}\\ i\sinh(2\lambda)e^{-ik\omega t}&\cosh(2\lambda)-ck/12\end{pmatrix}$. (108)
Note that the fundamental representation is nonunitary since $\mathrm{SU}$(1,1) is noncompact—as is manifest in the non-Hermitian matrix (108). However, the fundamental representation is faithful, so any lesson learned in this representation holds for $\mathfrak{su}(1,1)$ in general.
@@
By the definition of time-ordered exponentials as product integrals, the time-evolution operator (74) can be represented as in Eq. (75). The matrix representation (108) of the Hamiltonian thus yields
$U(t)\big|_{2\times 2}=e^{i\frac{\pi c}{12L}k^{2}t}\lim_{M\to\infty}\prod_{m=0,\ldots,M}^{\longleftarrow}\exp\biggl[\frac{\pi k}{L}\delta t\begin{pmatrix}i\cosh(2\lambda)&\sinh(2\lambda)e^{ik\omega m\delta t}\\ \sinh(2\lambda)e^{-ik\omega m\delta t}&-i\cosh(2\lambda)\end{pmatrix}\biggr$], (109)
@@
$\begin{pmatrix}0&e^{-i\omega s/2}\\ e^{i\omega s/2}&0\end{pmatrix}\begin{pmatrix}-i\bar{\alpha}&\bar{\beta}\\ \beta&i\alpha\end{pmatrix}\begin{pmatrix}0&e^{-i\omega s/2}\\ e^{i\omega s/2}&0\end{pmatrix}=\begin{pmatrix}i\alpha&\beta\,e^{-i\omega s}\\ \bar{\beta}\,e^{i\omega t}&-i\bar{\alpha}\end{pmatrix}$ (110)
for $s=-km\delta t$ and $\alpha=\cosh(2\lambda)$, $\beta=\sinh(2\lambda)$. It follows that each pair of consecutive factors in the product (109) gives rise to products of the form
$\begin{pmatrix}0&e^{ik\omega(m+1)\delta t/2}\\ e^{-ik\omega(m+1)\delta t/2}&0\end{pmatrix}\begin{pmatrix}0&e^{ik\omega m\delta t/2}\\ e^{-ik\omega m\delta t/2}&0\end{pmatrix}=\begin{pmatrix}e^{ik\omega\delta t/2}&0\\ 0&e^{ik\omega\delta t/2}\end{pmatrix}$, (111)
which are crucially independent of $m$. Thanks to this simplification, the matrix representation (109) of the evolution operator can be recast as
$U(t)\big|_{2\times 2}=e^{i\frac{\pi c}{12L}k^{2}t}\lim_{M\to\infty}\begin{pmatrix}0&e^{ik\omega t/2}\\ e^{-ik\omega t/2}&0\end{pmatrix}\!\mathsf{A}^{M}\exp\biggl[\frac{\pi k}{L}\delta t\begin{pmatrix}-i\cosh(2\lambda)&\sinh(2\lambda)\\ \sinh(2\lambda)&i\cosh(2\lambda)\end{pmatrix}\biggr]\!\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, (112)
@@
$\mathsf{A}=\exp\biggl[\frac{\pi k}{L}\delta t\begin{pmatrix}-i\cosh(2\lambda)&\sinh(2\lambda)\\ \sinh(2\lambda)&i\cosh(2\lambda)\end{pmatrix}\biggr]\begin{pmatrix}e^{ik\omega\delta t/2}&0\\ 0&-e^{ik\omega\delta t/2}\end{pmatrix}=\begin{pmatrix}a&b\\ \bar{b}&\bar{a}\end{pmatrix}+O(\delta t^{2})$ (113)
with $a=\bigl[1-i(\pi k/L)\delta t\cosh(2\lambda)\bigr]e^{ik\omega\delta t/2}$ and $b=(\pi k/L)\delta t\sinh(2\lambda)e^{-ik\omega\delta t/2}$. One can now diagonalize $\mathsf{A}$, noting that all $O(\delta t^{2})$ contributions give rise to subleading corrections in $1/M$, to find
$U(t)\big|_{2\times 2}$ $=e^{i\frac{\pi c}{12L}k^{2}t}\begin{pmatrix}e^{ik\omega t/2}&0\\ 0&e^{-ik\omega t/2}\end{pmatrix}$ (114)
@@
Lastly, we use the ansatz $e^{i[G+\bar{F}L_{-k}+EL_{0}+FL_{k}]}\big|_{2\times 2}$ ($G,E\in\mathbb{R}$ and $F\in\mathbb{C}$) for the second matrix in Eq. (114). The fundamental representation (107) then allows one to show that Eq. (76) holds for the evolution operator. This can also be used to find the diffeomorphism $g_{t}\in\operatorname{Diff}(S^{1})$ corresponding to the two last exponentials in Eq. (76). Owing to the properties of one-parameter flows, one can indeed show that $U(t)=e^{i\alpha_{t}}\,{\cal U}(g_{t})\,{\cal U}(g_{0})^{-1}$ in terms of a diffeomorphism $g_{t}$ as in Eq. (56), with $\Omega$ given by (78) and $\zeta(\varphi)$ a transformation (71) with $\alpha$, $\beta$ as in Eq. (77).
@@
When the initial condition is $g_{0}=f_{0}$ so that $|\psi(0)\rangle=\,{\cal U}(f_{0})|h\rangle$, one also needs to express ${\cal U}(f_{0})$ as a $2\times 2$ matrix for $f_{0}$ of the $\mathrm{SL}(2,\mathbb{R})$ form (71) with $\alpha=\cosh(\lambda)$, $\beta=\sinh(\lambda)$. This $f_{0}$ is the one-parameter flow generated by $(2/k)\sin(k\varphi)\partial_{\varphi}$ with respect to $\lambda$, so one finds
$\,{\cal U}(f_{0})\big|_{2\times 2}$ $=\exp\biggl(\frac{\lambda}{k}[L_{-k}-L_{k}]\biggr)\big|_{2\times 2}$ (115)
$=\begin{pmatrix}\cosh(\lambda)&i\sinh(\lambda)\\ -i\sinh(\lambda)&\cosh(\lambda)\\ \end{pmatrix}$.
Replicating the steps that led to Eq. (76), the above can be used to factorize the product of ${\cal U}(f_{0})$ with the evolution operator $U(t)$ as
$U(t)\,{\cal U}(f_{0})=e^{i\rho_{t}}e^{\chi_{t}L_{-k}-\bar{\chi}_{t}L_{k}}e^{i\sigma_{t}L_{0}}.$ (116)
Here, $\rho_{t}$ and $\sigma_{t}$ are real coefficients given by
$\rho_{t}$ $=\frac{\pi c}{12L}\Bigl[k^{2}t+\frac{L\sigma_{t}}{2\pi}\bigl(k^{2}-1\bigr)\Bigr],$ (117)
$\sigma_{t}$ $=-\omega t-\frac{2}{k}\arctan\Bigl(\Bigl[\frac{2\pi}{L\Omega}-\frac{\omega}{\Omega}\Bigr]\tan(k\Omega t/2)\Bigr),$
and $\chi_{t}$ is determined by $\eta_{t}$ in Eq. (80), with $\Omega$ given in Eq. (78). The ray of the vector $U(t)\,{\cal U}(f_{0})|h\rangle$ is only sensitive to $\chi_{t},\bar{\chi}_{t}$ as in Eq. (79), since $\rho_{t}$, $\sigma_{t}$ only contribute an overall phase $\rho_{t}+h\sigma_{t}$ due to Eq. (22).
Lastly, Eqs. (114) and (115) can be combined to represent ${\cal U}(f_{0})^{-1}U(t)\,{\cal U}(f_{0})$ as an $\mathrm{SU}(1,1)$ matrix:
${\cal U}(f_{0})^{-1}U(t)\,{\cal U}(f_{0})\big|_{2\times 2}=\begin{pmatrix}u_{t}&w_{t}\\ \bar{w_{t}}&\bar{u_{t}}\end{pmatrix}$ (118)
with $u_{t}$ as in Eq. (84) and
$w_{t}=\sinh(2\lambda)\biggl[$ $\frac{\omega}{\Omega}\sin(k\Omega t/2)\cos(k\omega t/2)$ (119)
$\begin{aligned} &-\cos(k\Omega t/2)\sin(k\omega t/2)\\ &+i\frac{2\pi}{L\Omega}\sin(k\Omega t/2)\sin(k\omega t/2)\biggr].\end{aligned}$
This must, on general grounds, be the nonunitary $2\times 2$-matrix representation of an element of the form $e^{D}e^{AL_{-k}}e^{BL_{0}}e^{CL_{k}}$. One finds that
$\begin{gathered}A=\frac{i}{k}\frac{\bar{w}}{u},\quad\;B=-\frac{2}{k}\log(u),\quad\;C=\frac{i}{k}\frac{w}{u},\\ e^{D}=e^{i\frac{\pi c}{12L}k^{2}t}e^{B\frac{c}{24}(k^{2}-1)}.\end{gathered}$ (120)
It follows using Eq. (22) that
$\langle\psi(0)|\psi(t)\rangle$ $=\langle h|\,{\cal U}(f_{0})^{-1}U(t)\,{\cal U}(f_{0})|h\rangle$ (121)
$=e^{D}\langle h|e^{BL_{0}}|h\rangle=e^{D+Bh},$
which yields Eq. (83) for the return amplitude.
@@ Appendix E Lattice calculations
Lattice Hamiltonian and adjacency matrix.
We consider a driven inhomogeneous free-fermion chain at half filling. The operator algebra is generated by fermionic creation (annihilation) operators $c_{j}^{\dagger}$ ($c_{j}$) at sites $j=1,\ldots,N$, which satisfy the usual anticommutation relations $\{c_{j^{\vphantom{\prime}}}^{\vphantom{\dagger}},c_{j^{\prime}}^{\dagger}\}=\delta_{j,j^{\prime}}$ and $\{c_{j^{\vphantom{\prime}}}^{\vphantom{\dagger}},c_{j^{\prime}}^{\vphantom{\dagger}}\}=\{c_{j^{\vphantom{\prime}}}^{\dagger},c_{j^{\prime}}^{\dagger}\}=0$. The Hamiltonian is
$H_{N}(t)=-\frac{1}{2}\sum_{j=1}^{N}J_{j}^{\vphantom{\dagger}}(t)\Bigl(c_{j}^{\dagger}c_{j+1}^{\vphantom{\dagger}}+c_{j+1}^{\dagger}c_{j}^{\vphantom{\dagger}}\Bigr)$ (122)
with site- and time-dependent nearest-neighbor hoppings $J_{j}(t)=J_{j}(t+T)>0$ and periodic boundary conditions, $c_{N+1}^{(\dagger)}=c_{1}^{(\dagger)}$ and $J_{N+1}^{\vphantom{\dagger}}(t)=J_{1}^{\vphantom{\dagger}}(t)$. A Jordan-Wigner transformation maps this model on the inhomogeneous spin chain (2) with $\Delta=0$ (up to boundary terms) [107]. We assume that the system size $L$ is much larger than the lattice spacing $L/N$, i.e. $N\gg 1$, and that $J_{j}(t)$ vary on mesoscopic length scales. Then, the model (122) can be effectively described in the low-energy regime by an inhomogeneous CFT (65) with $c=\bar{c}=1$, involving both right- and left-moving (chiral and antichiral) components [35].
The Hamiltonian for any periodic fermionic chain with nearest-neighbor hoppings can be represented by a periodic Jacobi matrix. There is no external field in our case, so the matrix is off-diagonal and reads
$\mathsf{J}(t)\equiv\begin{pmatrix}0&J_{1}(t)&0&\ldots&0&J_{N}(t)\\ J_{1}(t)&0&J_{2}(t)&\ldots&0&0\\ 0&J_{2}(t)&0&\ldots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&0&J_{N-1}(t)\\ J_{N}(t)&0&0&\ldots&J_{N-1}(t)&0\\ \end{pmatrix}$, (123)
which is nothing but the adjacency matrix corresponding to the Hamiltonian (122). The latter can be recast as $H_{N}(t)=(c_{1}^{\dagger},c_{2}^{\dagger},\ldots,c_{N}^{\dagger})\mathsf{H}(t)(c_{1}^{\vphantom{\dagger}},c_{2}^{\vphantom{\dagger}},\ldots c_{N}^{\vphantom{\dagger}})^{T}$ with $\mathsf{H}(t)\equiv-\mathsf{J}(t)/2$ interpreted as a single-particle Hamiltonian. If $\mathsf{U}$ denotes the unitary $N\times N$ matrix that diagonalizes the initial matrix Hamiltonian $\mathsf{H}(0)$, then one can write
$H_{N}(0)=\sum_{j=1}^{N}E_{j}(0)\gamma_{j}^{\dagger}\gamma_{j}^{\vphantom{\dagger}},\quad\;\gamma_{j}\equiv\sum_{\ell=1}^{N}(\mathsf{U}^{\dagger})_{j\ell}c_{\ell},$ (124)
where $E_{j}(0)$ denotes the eigenenergies of $\mathsf{H}(0)$ and $\gamma_{j}^{\dagger}$ ($\gamma_{j}^{\vphantom{\dagger}}$) are new fermionic creation (annihilation) operators. General properties of periodic Jacobi matrices [108] imply that the $E_{j}(0)$ can be ordered to be (nonstrictly) increasing in $j=1,\ldots,N$, with each eigenenergy equal to at most one other.
The many-body ground state is the filled Dirac sea, which corresponds to filling all single-particle states with negative energies. For simplicity, let $N$ be even. Assume furthermore that $N/2$ is odd, since that excludes that zero is an eigenenergy [108]. Then, at half filling, all states with $j=1,\ldots,N/2$ have negative energies, [^19] For $N$ even, let $S=\operatorname{diag}(1,-1,1,-1,\ldots,1,-1)$. Since $\mathsf{J}(t)$ in Eq. (123) satisfies $\mathsf{S}\mathsf{J}(t)=-\mathsf{J}(t)\mathsf{S}$, if $\mathbf{x}$ is an eigenvector of $\mathsf{H}(t)=-\mathsf{J}(t)/2$ with eigenvalue $E$, then $\mathsf{S}\mathbf{x}$ is also an eigenvector with eigenvalue E-E. i.e. the ground state is
$|\psi_{N}(0)\rangle=\gamma_{1}^{\dagger}\cdots\gamma_{N/2}^{\dagger}|\mathrm{vac}\rangle,$ (125)
where $|\mathrm{vac}\rangle$ denotes the state with no fermions. [For $N$ odd, one must replace $N/2$ by the maximal value $N_{\text{max}}$ for which $E_{N_{\text{max}}}(0)$ is negative.] Note that general filling is obtained by including the term $\sum_{j=1}^{N}\mu_{j}(t)(c_{j}^{\dagger}c_{j}^{\vphantom{\dagger}}-1/2)$ in the Hamiltonian (122), with a time-dependent inhomogeneous chemical potential $\mu_{j}(t)$. One then needs to fill its corresponding negative-energy one-particle states. These are states of the one-particle Hamiltonian given by Eq. (123), now with nonzero diagonal elements $\mu_{1}(t),\ldots,\mu_{N}$(t), which must have the same mesoscopic spatial dependence as $J_{j}(t)$ for there to be an emergent CFT at low energies (see Ch. 5 in [107]).
@@
Our considerations involve only fermions at zero temperature, so the (nonchiral) CFT model (65) is expected to be a good low-energy description of the lattice dynamics. Concretely, we let $|\psi_{N}(0)\rangle$ be the ground state (125) of $H_{N}(0)$, which in CFT language corresponds to the vacuum $|h\rangle\otimes|\bar{h}\rangle=|0\rangle\otimes|0\rangle$ as defined below Eq. (22). We then calculate the return probability after one period $T$,